# VEnvision3D: A Synthetic Perception Dataset for 3D Multi-Task Model Research

### 29 Feb 2024 v1

@@ Abstract
Developing a unified multi-task foundation model has become a critical challenge in computer vision research. In the current field of 3D computer vision, most datasets solely focus on a relatively limited set of tasks, which complicates the concurrent training requirements of various downstream tasks. This makes the training of multi-objective networks difficult to proceed with, which further hinders the development of foundation models in the 3D vision field. In this paper, we introduce VEnvision3D, a large 3D synthetic perception dataset for multi-task learning, including depth completion, segmentation, upsampling, place recognition, and 3D reconstruction. Since the data for each task was collected in the same scenarios, tasks are inherently aligned in terms of the utilized data. Therefore, such a unique attribute can assist in exploring the potential for the multi-task model and even the foundation model without separate training methods. Several new benchmarks based on the characteristics of the proposed dataset were presented. Extensive studies were performed on end-to-end models, revealing new observations, challenges, and opportunities for future research. In addition, we designed a straightfoward multi-task network to uncover the ability that VEnvision3D can offer for the foundation model. Our dataset and code will be open-sourced upon acceptance.
@@ 1 Introduction
The development of AI models has made many novel tasks possible. Previously, it was inconceivable that AI could autonomously imitate human creativity; however, VoiceBox [33] and DALLE-2 [50] are now capable of independently generating audio and images. Recently, foundation models (trained on a vast quantity of data at scale such that they can be adapted to a wide range of downstream tasks)[6] have once again sparked transformative revolutions across diverse industries. Examples include CLIP [49] and SAM [31] in computer vision, GPT-4 [42] in natural language processing, as well as UniAD [27] in autonomous driving. Multi-task, unified, end-to-end frameworks have become the new direction for the foundation model.
@@
The development of these models has benefited from the improvement of computing power, algorithm optimization, and the emergence of large-scale, high-quality data [73], such as the widely successful ImageNet [12] and InstructGPT [43]. Thanks to the availability of large-scale 2D image datasets and relatively small domain shifts between images [72], the development of the foundation model in 2D computer vision has become relatively mature. However, due to a relatively late start, we lack a large number of available datasets in 3D computer vision. As shown in Fig. 1, with a significantly smaller quantity of datasets compared to datasets in the 2D domain, some mainstream datasets also have much smaller scales [12, 37, 38, 17, 9, 4].
Figure 1: A comparison between dataset quantities in 2D and 3D vision and the data volume of classic datasets across different domains. The statistic data is from the Papers With Code[1] website.
@@
In the case of insufficient datasets, different parameter designs of 3D sensors result in a domain gap in 3D data that is much larger than that in 2D images, which further complicates the integration of the 3D data. Currently, a singular dataset is typically limited to specific downstream tasks. For example, the SemanticKITTI [4] dataset provides large-scale annotated point clouds, but the fixed collection route makes it unable to be used for place recognition tasks. Such a situation forces us to train multiple networks separately on different datasets, which not only affects the efficiency and accuracy of multi-task models but also limits the exploration of inter-task relationships [61]. Multi-task learning is considered the earliest attempt of the foundation model [16, 55]. However, the current situation of insufficient quantity of 3D datasets and the difficulty of data integration makes the research on 3D multi-task learning extremely challenging. Subsequently, this further impedes the development of foundation models in the 3D vision domain.
@@
A dataset collected within the same scenarios is required for investigating the relationships among subtasks, which can offer fresh insights in the development of 3D vision foundation models. In this paper, we introduce the VEnvision3D dataset, facilitating the exploration of multi-task models and the relationships between tasks in the field of 3D Vision, including depth completion, segmentation, upsampling, place recognition, and 3D reconstruction. Collected in CARLA Simulator [15], we divide the dataset into several subset by specific task. Meanwhile, based on the characteristics of virtual environments, the dataset provides functionalities that real datasets cannot provide, such as simulating temporal changes in the environment and sampling point clouds on model surfaces. This expands the range of scenarios and variation factors for evaluating algorithm performance, enabling more comprehensive exploration of their robustness and adaptability.
@@
Additionally, We conducted experiments to investigate the mutually beneficial relationships among specific subtasks. The experimental results revealed that this mutually beneficial relationship can improve the accuracy of sub-tasks. Beyond the scope of multi-task model development, we believe that a versatile dataset can assist researchers in exploring the potential for unified foundation models. Thus, we can offer a platform that fosters innovation for researchers(See Fig. 2), unifies the target tasks within the field of 3D Vision, and ultimately achieve a Foundation Model of Future-oriented Artificial General Intelligence.
@@
  • We introduce the VEnvision3D dataset, the first synthetic dataset that allows validation of multiple tasks within the same scenario.
  • The data collected in virtual environments yields novel attributes that are difficult to acquire in real-world situations, thereby offering fresh insights.
  • Using our dataset, we designed a straightforward multi-task network designed to demonstrate the mutual benefits among different tasks, highlighting the dataset’s versatility and potential to advance foundation model research.
@@ 2 Related Works
Dataset Cities Task
Dep. Ups. Seg. Recg. Recs.
Real-world KITTI [17] 1
SemanticKITTI [4] 1
ShapeNet [9] object
Oxford RobotCar [40] 1
S3DIS [3] indoor
Synthetic SynthCity [18] 1
STPLS3D [10] aerial
KITTI-CARLA [13] 8
AIODrive [64] 8
SHIFT [56] 8
Ours 8
Table 1: Comparison of supported tasks of existing 3D point cloud datasets. VEnvision3D is the only dataset that simultaneously supports depth completion(Dep.), segmentation(Seg.), upsample(Ups.), place recognition(Recg.), and 3D reconstruction(Recs.) in the same environmental domain.
Figure 2: The distinctive features of our dataset. We have innovated in five key functional areas.
Numerous datasets have been proposed in the past. However, they often exhibit limitations by being restricted to a single functionality. Tab. 1 shows the functionalities provided by current mainstream datasets. The following chapter presents a compilation of several renowned datasets.
Depth Completion. KITTI DC [58] is a widely used large-scale outdoor dataset that contains over 93,000 semi-dense depth maps with the corresponding raw sparse LiDAR scans and RGB images  [23]. The DenseLivox  [70] dataset is collected with much denser depth maps than KITTI. DeepLiDAR  [48] provides a synthetic dataset generated by the CARLA Simulator  [15]. Our dataset provides accurate ground truth depth obtained through rendering and point cloud projections with varying densities in the same frame.
Upsample. There are various point cloud datasets [65, 9, 60, 5] used to evaluate the performance of point cloud upsampling algorithms. Researchers usually downsample the ground truth as the input. Some researchers choose to build their dataset for training and testing of point cloud upsampling models  [69, 46, 68, 34, 47]. Compared to ours, these datasets do not provide data collected in large-scale outdoor scenes.
@@
3D Reconstruction. LiDAR-based SLAM has become a widely adopted method for 3D reconstruction. KITTI [17] is now the most popular benchmark for odometry evaluation. The NCLT dataset [8] comprises 27 extended sequences captured by Segway platforms. The UrbanNav dataset [21] offers a challenging data source for advancing the study of accurate and robust positioning in urban canyons. Our dataset includes city-level mesh models, enabling the development of novel benchmark methods.
3D Multi-task Dataset. Recently, some synthetic datasets with multiple sensors have been released. KITTI-CARLA [13] utilizes sensors identical to those in the KITTI dataset within a virtual environment. AIODrive [64] is generated with a specific focus on high-density long-range LiDAR datasets. SHIFT [56] features a comprehensive sensor suite and annotations, enabling the investigation of performance degradation in perception systems as domain shift increases. These datasets are collected exclusively through CARLA. However, they meet the multi-task requirements solely by enhancing sensor configurations, lacking organization based on specific downstream tasks, or exploring relationships between different tasks.
@@
16/32/64/128-channal LiDAR projected image; Day&Night monocular RGB image; rendered depth ground truth.
Upsample -
16/32/64/128-channal LiDAR scan.
@@
32/64/128-channal LiDAR scan with semantic&instance label. Provide semantic-labeled point clouds obtained from the mesh model.
@@
Monocular RGB image; rendered depth image; 64-channal LiDAR scan with pose and location. Four partially overlapping scanning routes collected before and after environmental migration.
@@
Stereo RGB image; rendered depth image; 64-channal LiDAR scan with pose and location. Provide city-level mesh models as ground truth.
Table 2: We organize the format of our dataset according to mainstream standards. This table contains the data items included in each task.
3 VEnvision3D Dataset
We have designed a data collection plan based on the requirements of mainstream point cloud tasks in the field. A comprehensive explanation of the dataset can be found in references Tab. 2 and Tab. 3.
@@ 3.1 Data Collection and Annotation
To make our data more user-friendly, we have adopted some of the mainstream dataset formats. We employed a LiDAR sensor, RGB Camera, Depth Camera, and GNSS Sensor for data collection purposes. To enhance the realism of the LiDAR point clouds, two augmentation methods are employed. Firstly, points are randomly removed based on their intensity values, with a higher probability of elimination for those with lower intensities. Secondly, a selection of points is randomly perturbed along the laser ray’s direction, introducing noisy distance measurements. In response to contemporary point cloud processing research, the dataset is categorized based on various functionalities, including depth completion, segmentation, upsampling, place recognition, and 3D reconstruction. This approach facilitates ease of use and enables the exploration of connections between different tasks. Thanks to the advantages of virtual environments, data annotation can be automatically performed. We provide instance and semantic segmentation labels on 28 classes and the real depth map generated by the renderer. We also obtain semantic-labeled point clouds through Monte Carlo sampling [20] for the simulated city model, providing a new benchmark for scene-level completion, 3D reconstruction, and semantic segmentation tasks.
3.2 Unique Feature
Different density data in the same frame. In real-world scenarios, it is challenging to obtain data with varying point densities within a single frame due to the limited availability of LiDAR sensors in vehicles. However, in our virtual environment, we have strategically placed LiDAR sensors with different channel counts (16/32/64/128 channels) at the same location. This unique configuration enables us to deliver, for the first time, scene-level point clouds with multiple densities in an upsampling task. The exploration of scene-level upsampling methods is crucial for the future development of intelligent point cloud processing. Additionally, we provide point clouds with varying densities for both depth completion and semantic segmentation tasks. These can be used to assess the algorithm’s robustness to density changes and for domain adaptation studies.
@@ 3.2 Unique Feature
16/32/64/128-channals, $360^{\circ}$ horizontal FoV, $-25^{\circ}$ to $30^{\circ}$ vertical FoV, 20Hz frequency, $\leq$ 100m range.
@@
Resolution of 1216 × 352, FoV of $90^{\circ}$, Aperture of $f/1.4$.
@@
GNSS
20Hz frequency.
Table 3: Sensor parameter settings.
Surface sampling from city-level model. The existing scene-level data is obtained through LiDAR measurements, inevitably leading to the occurrence of obstructions during the collection process, making it challenging to capture the complete scene. To overcome this issue, we have employed Monte Carlo Sampling to obtain point clouds and labels from the city mesh model’s surface in our dataset. Subsequently, we have divided the entire point cloud into blocks of 100m × 100m (consistent with the LiDAR detection distance), with a cropping interval of 6m. This approach ensures that the blocks simulate an unobstructed effect. By adopting this method, we aim to facilitate the exploration of techniques for scene-level completion of point clouds, while previous research primarily concentrated on completing individual objects. We also believe that a city-level model can introduce innovative evaluation methods for city-scale 3D reconstruction.
Simulate migration changes in environments. Traditional datasets are typically obtained through a one-time acquisition process, which fails to account for temporal environmental variations such as the seasonal decay of vegetation and the evolution of construction projects. Environmental changes are crucial for the development of adaptive algorithms, yet acquiring longitudinal real-world datasets that encompass these changes is a formidable challenge, often necessitating years of data collection. To address this gap, we have employed a virtual environment to simulate the construction of buildings by selectively editing existing structures. We designed four routes, including repetitive paths and closures. Vehicles traversed these routes before and after the simulated environmental transformations, allowing us to amass datasets reflecting different environmental parameters. We envisage that such an attribute can empower researchers to more thoroughly investigate the influence of environmental changes on place recognition algorithms.

### 05 Mar 2024 v2

@@ Abstract
Developing a unified multi-task foundation model has become a critical challenge in computer vision research. In the current field of 3D computer vision, most datasets only focus on single task, which complicates the concurrent training requirements of various downstream tasks. In this paper, we introduce VEnvision3D, a large 3D synthetic perception dataset for multi-task learning, including depth completion, segmentation, upsampling, place recognition, and 3D reconstruction. Since the data for each task is collected in the same environmental domain, sub-tasks are inherently aligned in terms of the utilized data. Therefore, such a unique attribute can assist in exploring the potential for the multi-task model and even the foundation model without separate training methods. Meanwhile, capitalizing on the advantage of virtual environments being freely editable, we implement some novel settings such as simulating temporal changes in the environment and sampling point clouds on model surfaces. These characteristics enable us to present several new benchmarks. We also perform extensive studies on multi-task end-to-end models, revealing new observations, challenges, and opportunities for future research. Our dataset and code will be open-sourced upon acceptance.
Keywords:
3D computer vision Datasets and evaluation Point cloud understanding Multi-task learning Environment simulation
@@ 1 Introduction
The development of AI models has made many novel tasks possible. In recent years, vision foundation models [6] (trained on a vast quantity of data at scale such that they can be adapted to a wide range of downstream tasks) have once again sparked transformative revolutions across diverse industries. Examples include CLIP [46] and SAM [31] in 2D computer vision, as well as UniAD [27] in autonomous driving. Multi-task, unified, end-to-end frameworks have become the new direction for the vision foundation model.
@@
The development of these models has benefited from the emergence of large-scale, high-quality data [69], such as the widely successful ImageNet [12]. Thanks to the availability of large-scale 2D image datasets and relatively small domain shifts between images [68], the development of the foundation model in 2D computer vision has become relatively mature. However, due to a relatively late start, we lack a large number of available datasets in 3D computer vision. As shown in Fig. 1, with a significantly smaller quantity of datasets compared to datasets in the 2D domain, some mainstream datasets also have much smaller scales [12, 36, 37, 17, 9, 4].
Refer to caption Figure 1: A comparison between dataset quantities in 2D and 3D vision and the data volume of classic datasets. The statistic data is from the Papers With Code[1] website. Dataset Cities Task Dep. Ups. Seg. Recg. Recs. Real-world KITTI [17] 1 SemanticKITTI [4] 1 ShapeNet [9] object Oxford RobotCar [39] 1 S3DIS [3] indoor Synthetic SynthCity [18] 1 STPLS3D [10] aerial KITTI-CARLA [13] 8 AIODrive [60] 8 SHIFT [52] 8 VEnvision3D (Ours) 11 Table 1: Comparison of supported tasks of existing 3D point cloud datasets. VEnvision3D is the only dataset that simultaneously supports depth completion(Dep.), segmentation(Seg.), upsampling(Ups.), place recognition(Recg.), and 3D reconstruction(Recs.) in the same environmental domain.
@@
With only a handful of datasets available to 3D vision researchers, parameter differences between 3D sensors lead to larger domain gap than 2D images, making the situation even more unpromising. This large gap further allows a dataset to be used for only one task for which it was designed. For example, the SemanticKITTI [4] dataset provides large-scale annotated point clouds for segmentation, but the fixed collection route makes it unable to be used for place recognition task. Such a situation forces us to train multiple networks separately on different datasets, which not only affects the efficiency and accuracy of multi-task models but also limits the exploration of inter-task relationships [57].
@@
A dataset collected within the same environmental domain is required for investigating the relationships among sub-tasks, which can offer fresh insights into the development of 3D vision multi-task models. In this paper, we introduce VEnvision3D dataset, facilitating the exploration of multi-task models and the relationships between tasks in the field of 3D Vision, including depth completion, segmentation, upsampling, place recognition, and 3D reconstruction. Using CARLA Simulator [15], we divide the dataset into several subsets by specific task. To unify the data collection of multiple sub-tasks, it necessitates a meticulously designed distribution of sensors, a setup that presents more challenges compared to collecting single data types. The advantage of virtual environments being freely editable makes this possible. Concurrently, in the process of exploration, we also implement features that were previously difficult to realize in the real world, such as simulating temporal changes in the environment and sampling point clouds on model surfaces. We manifest data, which in real life incurs high costs, within virtual worlds, thereby providing preliminary data support for more cutting-edge research in the future. This expands the range of scenarios and variation factors for evaluating algorithm performance, enabling a more comprehensive exploration of their robustness and adaptability.
@@
Additionally, we conduct experiments to investigate the mutually beneficial relationships among specific tasks. To explore this initially, we design an multi-task end-to-end framework to leverage the features between different sub-tasks, thereby achieving semantic segmentation and upsampling concurrently. The experimental results tentatively reveal that this mutually beneficial relationship can improve the accuracy of sub-tasks. Beyond the scope of multi-task model development, we believe that our dataset can assist researchers in exploring the potential for unified models. Thus, we can offer a platform that fosters innovation for researchers (See Fig. 2), unify the target tasks within the field of 3D Vision, and ultimately achieve a foundation model of future-oriented artificial general intelligence.
Figure 2: The distinctive features of VEnvision3D dataset. Taking advantage of freely editable virtual environments, we implement many novel settings, such as the stacking of sensors, the generation of annotated city mesh models, and the simulation of environmental changes. These attributes enable us to realize multi-task functionalities and various types of data that are challenging to achieve in the real world.
@@
  • We introduce the VEnvision3D dataset, the first synthetic dataset that allows validation of multiple tasks within the same environmental domain.
  • The data collected in virtual environments yields novel attributes that are difficult to acquire in real-world situations, thereby providing new challenging benchmarks.
  • Using our dataset, we design a multi-task end-to-end network to demonstrate the mutual benefits among two different tasks, highlighting the dataset’s versatility and potential to advance future research.
@@ 2 Related Works
Numerous 3D datasets have been proposed in the past. However, they often exhibit limitations by being restricted to a single functionality. Tab. 1 shows the functionalities provided by current mainstream datasets. We introduce the classic 3D perception datasets as follows:
2.1 Single-task Dataset
Depth Completion. Currently, datasets for depth completion tasks are relatively scarce. KITTI DC [54] is a widely used large-scale outdoor dataset that contains over 93,000 semi-dense depth maps with the corresponding raw sparse LiDAR scans and RGB images [23]. The DenseLivox [66] dataset is collected with much denser depth maps than KITTI. DeepLiDAR [45] provides a synthetic dataset generated by the CARLA Simulator [15]. In comparison, our dataset provides accurate ground truth depth obtained through rendering and point cloud projected image with varying densities in the same frame.
Upsampling. There are various point cloud datasets [61, 9, 56, 5] used to evaluate the performance of point cloud upsampling algorithms. For object-level upsampling tasks, researchers usually downsample the ground truth as the input. Some researchers choose to build their dataset for training and testing of point cloud upsampling models [65, 43, 64, 33, 44]. Compared to ours, these datasets do not provide data collected in large-scale outdoor scenes. For the first time, we have achieved point clouds of natively varying densities at the sensor level.
@@ 2.1 Single-task Dataset
3D Reconstruction. LiDAR-based SLAM has become a widely adopted method for 3D reconstruction. KITTI Odometry [17] is now the most popular benchmark for odometry evaluation. The NCLT dataset [8] comprises 27 extended sequences captured by Segway platforms. The UrbanNav dataset [21] offers a challenging data source for advancing the study of accurate and robust positioning in urban canyons. Our dataset includes city-level mesh models, enabling the development of novel benchmark methods.
While the aforementioned datasets have achieved commendable outcomes within their respective domains, it is imperative to acknowledge the profound discrepancies among 3D sensors (including brand, frequency, channels, and the number of emission points), which lead to significant domain shifts among these datasets. This necessitates individual training on each dataset, further complicating the integration of networks for various sub-tasks. There is a pressing need for a multi-task dataset that would enable researchers to design a multi-task framework in a singular effort.
2.2 Multi-task Dataset
Recently, some synthetic datasets with multiple sensors have been released. KITTI-CARLA [13] utilizes sensors identical to those in the KITTI dataset within a virtual environment. AIODrive [60] is generated with a specific focus on high-density and long-range LiDAR datasets. SHIFT [52] features a comprehensive sensor suite and annotations, enabling the investigation of performance degradation in perception systems as domain shift increases. These datasets are collected exclusively through CARLA.
However, these datasets merely utilize combinations of multiple sensors to fulfill multi-task requirements, with a predominant focus on 2D vision. They achieve multi-task capabilities by providing images annotated with different information. In 3D vision, sub-tasks are delineated with greater specificity and cannot be simply represented through images. This necessitates more meticulous design efforts to develop datasets capable of satisfying the requirements of multiple sub-tasks. Our dataset successfully provides the necessary data for a variety of 3D vision sub-tasks within the same environmental domain. Moreover, leveraging the advantages of virtual environments, we have collected many forms of data that are challenging to obtain in the real world, offering more solid support for 3D multi-task learning.
3 VEnvision3D Dataset
In order to make our dataset available for different sub-tasks in the same frame, we design a data acquisition plan capable of accommodating multiple sub-tasks. Specific data details are presented in Tab. 2, the sensor setup is illustrated in Fig. 3, and the parameters of the sensors are listed in Tab. 3. VEnvision3D has more than 140,000 fully annotated frames, occupies more than 2,500 GB, and has a total of 13 scene variations ranging from urban to rural. Different sub-tasks can share the same frame. We additionally provide city models and semantic-annotated point clouds sampled from their surfaces.
@@ 3 VEnvision3D Dataset
16/32/64/128-channel LiDAR projected image; monocular RGB image; rendered depth ground truth.
Upsampling -
16/32/64/128-channel LiDAR scan.
@@
16/32/64/128-channel LiDAR scan with semantic & instance label.
Provide semantic-annotated point clouds sampled from the city mesh model.
@@
Monocular RGB image; rendered depth image; 64-channel LiDAR scan with pose and location.
Four partially overlapping scanning routes collected before and after environmental migration.
@@
Stereo RGB image; 64-channel LiDAR scan with pose and location.
Provide city mesh models as ground truth.
Table 2: We organize the format of VEnvision3D according to mainstream standards. This table includes the data required for each sub-task. We can facilitate multi-task learning by providing data for different tasks in the same frame.
@@ 3.1 Data Collection and Annotation
We employ LiDAR with different channel count, RGB camera, Depth camera and GNSS/IMU for data collection purposes. To enhance the realism of the LiDAR point clouds, we manually add noise on point clouds. A selection of points is randomly perturbed along the laser ray’s direction, introducing noisy distance measurements. Taking advantage of the fact that virtual environments can be freely edited, we implement stacking of LiDAR sensors at the same location, generating point clouds that do not obscure each other and share exactly the same coordinate system. Thus we realize the collection of point clouds with different densities at the sensor level, which is the main reason why our dataset can help multi-task learning.
We allow the vehicle to follow random routes in the city map and simulate real traffic conditions in order to collect data in different scenarios. These city maps vary in size, with the largest covering 10,000 hectares and encompassing a range of scenes from urban to rural, and from streets to highways. The screenshots can be found in the supplementary materials. Note that for sub-tasks such as the 3D reconstruction and place recognition, which require fixed routes, we design several separate routes that can fulfill the task requirements. These routes overlap with the previously random ones and can also meet the requirements of other types of sub-tasks.
Figure 3: The vehicle system and the sensor layout. We successfully stack LiDARs with different channel count at the same location to simulate point clouds of different densities at the sensor level, and the data are automatically annotated. The point clouds generated by LiDAR sensors share the exact same coordinate system and do not produce mutual occlusions.
@@
16/32/64/128-channel, $360^{\circ}$ horizontal FoV, $-25^{\circ}$ to $30^{\circ}$ vertical FoV, 10Hz frequency, $\leq$ 120m range. With manually added noise.
@@
Resolution of 1216 × 352, FoV of $90^{\circ}$, Aperture of $f/1.4$. With post-process effects.
@@
GNSS/IMU
10Hz frequency.
Table 3: Sensor parameter settings. We meticulously configure our setup in accordance with the parameters of KITTI[17] sensors, and added noise and post-effects to ensure the realism of the data, aiming to minimize domain shift.
Thanks to the advantages of virtual environments, data annotation can be automatically performed. We provide instance and semantic segmentation labels on 28 classes and the real depth map generated by the renderer. We also obtain semantic-labeled point clouds through Monte Carlo sampling [20] from the city model, providing a new benchmark for scene-level completion, 3D reconstruction, and semantic segmentation tasks.
In response to contemporary point cloud processing research, the dataset is categorized based on various functionalities, including depth completion, segmentation, upsampling, place recognition, and 3D reconstruction. This approach facilitates ease of use and enables the exploration of connections between different tasks.
3.2 Unique Features
Different density data in the same frame. In real-world scenarios, it is challenging to obtain data with varying point densities within a single frame due to the limited availability of LiDAR in vehicles. However, in our virtual environment, we strategically place LiDAR with different channel counts (16/32/64/128 channels) at the same location. This unique configuration enables us to deliver, for the first time, scene-level point clouds with multiple densities in an upsampling task at sensor level. The exploration of scene-level upsampling methods is crucial for the future development of intelligent point cloud processing. Additionally, we provide point clouds with varying densities for both depth completion and semantic segmentation tasks. These can be used to assess the algorithm’s robustness to density changes and for domain adaptation studies.
Surface sampling from city-level model. The existing scene-level data is obtained through LiDAR measurements, inevitably leading to the occurrence of obstructions during the collection process, making it challenging to capture the complete scene. To overcome this issue, we employe Monte Carlo Sampling to obtain point clouds and labels from the city mesh model’s surface in our dataset. Subsequently, we divide the entire point cloud into blocks of 100m × 120m (consistent with the LiDAR detection distance), with a cropping interval of 6m. This approach ensures that the blocks simulate an unobstructed effect. By adopting this method, we aim to facilitate the exploration of techniques for scene-level completion of point clouds, while previous research primarily concentrated on completing individual objects. We also believe that a city-level model can introduce innovative evaluation methods for city-scale 3D reconstruction.
Simulate migration changes in environments. Traditional datasets are typically obtained through a one-time acquisition process, which fails to account for temporal environmental variations such as the seasonal decay of vegetation and the evolution of construction projects. Environmental changes are crucial for the development of adaptive algorithms, yet acquiring longitudinal real-world datasets that encompass these changes is a formidable challenge, often necessitating years of data collection. To address this gap, we use the virtual environment to simulate the construction of buildings by selectively editing existing structures. We design four routes, including repetitive paths and closures. Vehicles traverse these routes before and after the simulated environmental migration, allowing us to collect datasets reflecting different environmental parameters. We envisage that such an attribute can empower researchers to more thoroughly investigate the influence of environmental changes on place recognition algorithms.